OpenStudy (anonymous):

Vectors A and B have the same magnitude. Given that the magnitude of A + B is 54 times greater than the magnitude of A - B, find the angle between them?

OpenStudy (anonymous):

|A + B| = 54 |A -B| |A| + |B| + 2AB cos t = 54( |A| + |B| - 2AB cos t)

OpenStudy (anonymous):

A+B=sqrtA^+B^2+2ABCOSx